OpenStudy (anonymous):

Which expression is equivalent to x^5 • x^2? x^3 x^7 x^10 (x + x)^7

OpenStudy (cwrw238):

add the indices a^n * a^m = a^(n+m)

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

when you multiply terms like the above you add the exponents (indices) eg x^2 * x^8 = x ^(2+8) = x^10

OpenStudy (anonymous):

so x^7

OpenStudy (cwrw238):

right